A straight line through the origin meet The parallel line 4 x + 2 y = 9 and 2 X + Y + 6 = 20 at the point P and Q respectively then the point O divides the segment PQ in the ratio

Grade:11

1 Answers

Arun
6 years ago
 

The given parallel lines are 4x + 2y = 9 and 2x + y + 6 = 0.

The distance of the origin from 4x + 2y – 9 = 0 is |-9|/√42 + 2= 9/√20.

Similarly, the distance of origin from 2x+ y + 6 = 0 is

 |6|/√22 + 1= 6/√5.

Hence, the required ratio is (9/√20) / (6/√5) = ¾.